OpenStudy (anonymous):

The probabilty of winning a particular game at the game place is 1/4. Find the odds against winning the game. A. 1 to 4 B. 3 to 4 C. 3 to 1 D. 5 to 1

OpenStudy (anonymous):

B. 3 to 4 = 1 - 1/4 = 3/4